OpenStudy (anonymous):

A quality control inspector is in charge of inspecting drill bits that are produced on an assembly line. The diameter of a drill bit can be modeled by the equation |x − 0.525| = 0.01. Show your work in solving the equation to determine the minimum and maximum diameters of the drill bits.

OpenStudy (paki):

x − 0.525 = +0.01 and x − 0.525 = -0.01 solve for maximum and minimum.... x=0.01+0.525 and x= -0.01+0.525 x= 0.5625 and x= 0.515